# Let Androids Dream of Electric Sheep: A Human-like Image Implication Understanding and Reasoning Framework

### v1

@@ Abstract
Metaphorical comprehension in images remains a critical challenge for AI systems, as existing models struggle to grasp the nuanced cultural, emotional, and contextual implications embedded in visual content. While multimodal large language models (MLLMs) excel in basic Visual Question Answer (VQA) tasks, they struggle with a fundamental limitation on image implication tasks: contextual gaps that obscure the relationships between different visual elements and their abstract meanings. Inspired by the human cognitive process, we propose Let Androids Dream (LAD), a novel framework for image implication understanding and reasoning. LAD addresses contextual missing through the three-stage framework: (1) Perception: converting visual information into rich and multi-level textual representations, (2) Search: iteratively searching and integrating cross-domain knowledge to resolve ambiguity, and (3) Reasoning: generating context-alignment image implication via explicit reasoning. Our framework with the lightweight GPT-4o-mini model achieves SOTA performance compared to 15+ MLLMs on English image implication benchmark and a huge improvement on Chinese benchmark, performing comparable with the GPT-4o model on Multiple-Choice Question (MCQ) and outperforms 36.7% on Open-Style Question (OSQ). Additionally, our work provides new insights into how AI can more effectively interpret image implications, advancing the field of vision-language reasoning and human-AI interaction. Our project is publicly available at https://github.com/MING-ZCH/Let-Androids-Dream-of-Electric-Sheep.
@@ 1 Introduction
Metaphors are not just abstract concepts found in literature; they are also prevalent in our daily lives. For instance, when we say "time is money" or "life is a journey," we are using metaphors to convey complex ideas in a more contextual and understandable way. These metaphors highlight the integral role that metaphoric thinking plays in human communication and cognition. Just as we use metaphors to make sense of the world around us, we aim to enable AI to understand metaphors in a human-like manner. In linguistic terms, as George Lakoff and Mark Johnson elaborated in "Metaphors We Live By" [11], metaphors are not merely ornamental language devices but fundamental cognitive tools that allow us to conceptualize our surroundings. Metaphors possess characteristics such as systematicity, the creation of similarity, and imaginative rationality. Through cross-domain mapping, one concept can be used to comprehend another, allowing for a more insightful interpretation.
With the rapid advancement of large language models (LLMs), models such as OpenAI o1 [20], DeepSeek-R1 [4], and QwQ [25] have demonstrated remarkable text-reasoning capabilities. However, a significant amount of knowledge in the real world cannot be fully represented by text alone. Visual information, for instance, contains a wealth of knowledge that is not easily captured through text. As a result, there has been a growing interest in integrating visual information into text-reasoning tasks.
@@
Compared to language, vision is inherently complex, with its diverse representation of information, subjective understanding, and the difficulty in quantifying its information. In recent years, multimodal reasoning models such as QVQ [24] and K1.5 [23] have achieved outstanding performance. For example, K1.5 model has reached a high score on math, code and multimodal reasoning benchmarks [9, 14, 16, 26, 36]. However, these models still perform poorly on image metaphor questions [15, 37]. They tend to focus on the superficial elements of the image, neglecting the deeper connections and emotional expressions among these elements, as shown in Figure 1. It is important to note that these models excel at logical reasoning tasks, which are based on a different set of cognitive principles compared to image metaphor tasks. In contrast to the VQA task, which primarily centers on concrete image comprehension, the image metaphor entails a stronger emphasis on abstract meaning and higher-order reasoning capabilities. It is not a simple logical reasoning task and requires a different method to understand and generate implications. It requires the model to understand complex and abstract information, such as metaphors, symbols, and emotions in the image, rather than just the concrete contents.
@@
Image implication tasks consist of two main aspects: understanding and generation. Understanding image implication is a more complex and challenging task than understanding conventional images. It requires advanced cognitive abilities such as multi-hop reasoning and a sophisticated theory of mind (ToM), which are inherent to human cognition [15, 37]. Compared to understanding, generating implication is even more difficult. The fundamental challenge stems from the lack of contextual understanding of the key elements and internal relationships of the image. This lack of context hinders our ability to decipher the intended message or to create images that effectively convey specific meanings. Without the background of cultural, historical, or environmental context, the significance of key visual components remains elusive, impeding both interpretation and creative expression.
@@
Existing methods for solving the image metaphor understanding can be mainly divided into two categories: explicit metaphor mapping and model implicit reasoning. The former achieves image metaphor understanding by establishing a correspondence between metaphor ontology and visual representation. For example, the CLOT method [40] realizes image metaphor understanding through the mapping between metaphor ontology and visual representation. Model implicit reasoning relies on the model’s reasoning ability and does not require the explicit mapping construction. For example, C4MMD method [32] adopts an untrained chain-of-reasoning approach. However, explicit metaphor mapping, although it can provide a clear mapping, has limitations when dealing with complex many-to-many mappings and dynamically changing cultural backgrounds. On the other hand, model implicit reasoning, despite its potential, still faces challenges in handling complex metaphor understanding tasks, especially in situations involving multimodal information and cultural backgrounds.
@@
To address these problems, we analyze how humans understand metaphors and find that the essence of the difficulty in metaphor understanding and generation is contextual missing. Therefore, we propose a novel framework that more closely aligns with human cognitive processes for metaphor interpretation. Our framework first transforms visual information into textual representations and then iteratively searches to enrich these representations with out-of-domain knowledge, enabling deeper inferential reasoning. Experiments from both Multiple-Choice Question and Open-Style Question consistently verify the superiority of the proposed framework.
@@
  • We systematically analyze image implication tasks and find the difficulty of the metaphor understanding and reasoning task lies in contextual missing. From the perspective of human cognition, we proposed a new direction for solving these tasks – Contextual Alignment.
  • We propose a novel human-like three-stage framework Let Androids Dream (LAD), which implements the lightweight GPT-4o-mini model to achieve SOTA on English image implication benchmark and a huge improvement on Chinese image implication benchmark, comparable with the GPT-4o model and other top closed-source models on Multiple-Choice Question.
  • We design the challenging Open-Style Question (OSQ) with comprehensive metric to automatic evaluate the image implication tasks. This metric aligns 95.7% with human annotations, making it more suitable for diverse evaluation. Our LAD outperforms the GPT-4o model 36.7% on OSQ.
@@ 2.1 Image Implication
Image implication encompasses various cognitive aspects, including humor, sarcasm, and broader metaphorical understanding. Early research in this domain focused on specialized aspects, such as humor recognition [6, 7] and sarcasm detection [5]. As the rapid development of large language models (LLMs) brings new opportunities for analyzing image implication, we need more comprehensive evaluation frameworks. DeepEval [34] provided a systematic taxonomy of image implications. Subsequently, II-Bench [15] emerged as the first English image implication benchmark, followed by CII-Bench [37], which extended this evaluation framework to Chinese images. Image implication understanding requires sophisticated multi-hop reasoning and theory of mind (ToM) capabilities [15, 37]. Existing approaches fall into two categories: explicit metaphor mapping and model implicit reasoning. The first approach, represented by CLOT [40], constructs mappings between metaphor ontologies and visual representations. However, this approach faces key challenges: metaphorical relationships have complex many-to-many mappings that are difficult to formalize, and cultural references are too dynamic for static mappings. The second approach, exemplified by C4MMD [32], employs training-free CoT reasoning. Despite its promise, this approach struggles with the complex nature of metaphorical understanding, which surpasses traditional reasoning. The large search space for out-of-domain reasoning and changing cultural contexts limits its effectiveness. To address this, we propose a novel methodology that transforms visual information into texts and iteratively enriches them with out-of-domain knowledge, better aligning with human cognitive processes.
@@ 2.2 Vision-language Reasoning
The rapid advancement of LLMs has demonstrated remarkable text reasoning capabilities, as evidenced by models such as o1 [20], DeepSeek-R1 [4], and QwQ [25, 33]. However, real-world knowledge often transcends textual representation, with visual information encapsulating substantial world knowledge that pure language models cannot access. For example, images inherently contain rich, multi-layered information that often resists straightforward textual description, including spatial relationships, contextual nuances, and implicit knowledge that humans process intuitively. This limitation has driven research toward integrating visual information into text-based reasoning frameworks. Current research has developed three primary approaches to incorporate visual information into model reasoning: 1) Comprehensive MLLM Description: This approach treats visual content as a text grounding problem, as demonstrated by LLAVA-COT [31] and Mulberry [35]. 2) Multi-turn MLLM Interaction: Models like VoCoT [13] and V* [28] employ iterative question-answering to extract fine-grained visual information at various levels of detail. 3) Tool-augmented Reasoning: Frameworks such as Visual Sketchpad [8] and Whiteboard-of-Thought [17] leverage tool-based approaches to modify images and augment reasoning with prior knowledge embedded in these tools.
@@ 3.2 Stage II: Search
    WebSearch: For questions requiring external, dynamic, or highly specific information, WebSearch is invoked. Inspired by LLM search methods like MindSearch [3], but focusing on image implication tasks, our WebSearch component first employs the planner. The planner, acting as a high-level strategist, decomposes the initial search question into a series of more granular sub-questions. These sub-questions are structured into a directed acyclic graph (DAG), simulating a multi-step, exploratory information-seeking process. Subsequently, the searcher executes this plan. It performs hierarchical information retrieval for each sub-question from the internet, gathering relevant snippets and facts. This multi-agent method, with distinct planner and searcher modules, allows for parallel processing and dynamic refinement of the search strategy. The retrieved information for sub-questions is then synthesized to answer the original search question. This ensures access to recent developments and a broad spectrum of public knowledge, crucial for understanding contemporary image implications.
@@ 4.1 Baselines
Qwen2.5-VL-7B [2] 46% 40% 2.34 2.58
DeepSeek-VL2 [29] 46% 36% 2.82 2.86
Gemini-2.0-flash [22] 70% 68% 1.60 3.12
QwenVL-2.0-72B [27] 68% 54% 2.84 3.04
QwenVL-2.5-72B [2] 72% 56% 1.56 3.12
GLM-4V-plus [39] 64% 64% 3.01 3.12
Gemini-2.0-pro [22] 68% 62% 1.66 3.18
Grok-3 [30] 66% 64% 3.24 2.96
Claude-3.5-Sonnet [1] 68% 62% 3.22 3.78
GPT-4o [19] 74% 58% 2.94 3.76
GPT-4.1 [19] 74% 62% 3.30 3.92
@@
Gemini-2.0-flash-thinking [22] 64% 68% 1.66 2.84
QVQ-72B [24] 62% 56% 3.10 3.42
Doubao-1.5-thinking-vision-pro [21] 66% 66% 3.16 3.90
Grok-3-reasoning [30] 74% 64% 3.06 2.92
@@
Table 1: Overall results of different models on Multiple-Choice Question and Open-Style Question. The best-performing model in each category is in-bold, and the second best is underlined.
@@
Evaluation. Our evaluation utilizes two comprehensive image implication benchmarks, II-Bench [15] and CII-Bench [37], both featuring Multiple-Choice Question (MCQ). Furthermore, we manually construct a high-quality benchmark by randomly selecting 50 images from varied image types like illustrations and comics. And we measure accuracy by comparing the model’s selected option to the ground truth. Aware of potential MCQ biases [12, 18, 38] and the greater difficulty of generation over judgment tasks, we introduce a novel evaluation method Open-Style Question (OSQ). It uses the same images with the fixed question: “What is the implication in this image?”. And we use GPT-4o with a specialized evaluation metric as evaluators, validated by multiple human consistency checks.
Figure 3: A case study of different methods on Multiple-Choice Question. The End2End method shows superficial reasoning and the CoT method shows over-inference, while our LAD framework shows the correct path via more contextual alignment analysis. The full prompt is listed in Appendix D.
@@ 4.2.2 Results and Analysis
Table 1 presents comprehensive results of MCQ across different MLLMs on our high-level benchmark. The LAD framework demonstrates remarkable effectiveness, achieving SOTA performance with the lightweight GPT-4o-mini model. In English MCQ, our framework matches the performance of closed-sourced models like GPT-4o, GPT-4.1, and Grok-3-reasoning (74%), while significantly outperforming Claude-3.5-Sonnet and Gemini-2.0-pro by 9%. For Chinese MCQ, our framework achieves comparable results to GPT-4o, while substantially surpassing DeepSeek-VL2 by 44.4%.

### v2

# Let Androids Dream of Electric Sheep: A Human-Inspired Image Implication Understanding and Reasoning Framework

@@ Abstract
Metaphorical comprehension in images remains a critical challenge for AI systems, as existing models struggle to grasp the nuanced cultural, emotional, and contextual implications embedded in visual content. While multimodal large language models (MLLMs) excel in general Visual Question Answer (VQA) tasks, they struggle with a fundamental limitation on image implication tasks: contextual gaps that obscure the relationships between different visual elements and their abstract meanings. Inspired by the human cognitive process, we propose Let Androids Dream (LAD), a novel framework for image implication understanding and reasoning. LAD addresses contextual missing through the three-stage framework: (1) Perception: converting visual information into rich and multi-level textual representations, (2) Search: iteratively searching and integrating cross-domain knowledge to resolve ambiguity, and (3) Reasoning: generating context-alignment image implication via explicit reasoning. Our framework with the lightweight GPT-4o-mini model achieves SOTA performance compared to 15+ MLLMs on English image implication benchmark and a huge improvement on Chinese benchmark, performing comparable with the Gemini-3.0-pro model on Multiple-Choice Question (MCQ) and outperforms the GPT-4o model 36.7% on Open-Style Question (OSQ). Generalization experiments also show that our framework can effectively benefit general VQA and visual reasoning tasks. Additionally, our work provides new insights into how AI can more effectively interpret image implications, advancing the field of vision-language reasoning and human-AI interaction. Our project is publicly available at https://github.com/MING-ZCH/Let-Androids-Dream-of-Electric-Sheep.
@@ 1 Introduction
Metaphors are not just abstract concepts found in literature; they are also prevalent in our daily lives. For instance, when we say "time is money" or "life is a journey," we are using metaphors to convey complex ideas in a more contextual and understandable way. These metaphors highlight the integral role that metaphoric thinking plays in human communication and cognition. Just as we use metaphors to make sense of the world around us, we aim to enable AI to understand metaphors in a human-like manner. In linguistic terms, as George Lakoff and Mark Johnson elaborated in "Metaphors We Live By" [lakoff2008metaphors], metaphors are not merely ornamental language devices but fundamental cognitive tools that allow us to conceptualize our surroundings. Metaphors possess characteristics such as systematicity, the creation of similarity, and imaginative rationality. Through cross-domain mapping, one concept can be used to comprehend another, allowing for a more insightful interpretation.
With the rapid advancement of large language models (LLMs), models such as OpenAI o1 [o1], DeepSeek-R1 [deepseek-R1], and QwQ [qwq-32b-preview] have demonstrated remarkable text-reasoning capabilities. However, a significant amount of knowledge in the real world cannot be fully represented by text alone. Visual information, for instance, contains a wealth of knowledge that is not easily captured through text. As a result, there has been a growing interest in integrating visual information into text-reasoning tasks.
@@
Compared to language, vision is inherently complex, with its diverse representation of information, subjective understanding, and the difficulty in quantifying its information. In recent years, multimodal reasoning models such as QVQ [qvq-72b-preview] and K1.5 [team2025kimi] have achieved outstanding performance. For example, K1.5 model has reached a high score on math, code and multimodal reasoning benchmarks [jain2024livecodebench, lightman2023letsverifystepstep, lu2024mathvista, wang2024measuring, yue2023mmmu]. However, these models still perform poorly on image metaphor questions [liu2024iibenchimageimplicationunderstanding, zhang2024mllmsunderstanddeepimplication]. They tend to focus on the superficial elements of the image, neglecting the deeper connections and emotional expressions among these elements, as shown in Figure 1. It is important to note that these models excel at logical reasoning tasks, which are based on a different set of cognitive principles compared to image metaphor tasks. In contrast to the VQA task, which primarily centers on concrete image comprehension, the image metaphor entails a stronger emphasis on abstract meaning and higher-order reasoning capabilities. It is not a simple logical reasoning task and requires a different method to understand and generate implications. It requires the model to understand complex and abstract information, such as metaphors, symbols, and emotions in the image, rather than just the concrete contents.
@@
Image implication tasks consist of two main aspects: understanding and generation. Understanding image implication is a more complex and challenging task than understanding conventional images. It requires advanced cognitive abilities such as multi-hop reasoning and a sophisticated theory of mind (ToM), which are inherent to human cognition [liu2024iibenchimageimplicationunderstanding, zhang2024mllmsunderstanddeepimplication]. Compared to understanding, generating implication is even more difficult. The fundamental challenge stems from the lack of contextual understanding of the key elements and internal relationships of the image. This lack of context hinders our ability to decipher the intended message or to create images that effectively convey specific meanings. Without the background of cultural, historical, or environmental context, the significance of key visual components remains elusive, impeding both interpretation and creative expression.
@@
Existing methods for solving the image metaphor understanding can be mainly divided into two categories: explicit metaphor mapping and model implicit reasoning. The former achieves image metaphor understanding by establishing a correspondence between metaphor ontology and visual representation. For example, the CLOT method [zhong2024lets] realizes image metaphor understanding through the mapping between metaphor ontology and visual representation. Model implicit reasoning relies on the model’s reasoning ability and does not require the explicit mapping construction. For example, C4MMD method [xu-etal-2024-exploring] adopts an untrained chain-of-reasoning approach. However, explicit metaphor mapping, although it can provide a clear mapping, has limitations when dealing with complex many-to-many mappings and dynamically changing cultural backgrounds. On the other hand, model implicit reasoning, despite its potential, still faces challenges in handling complex metaphor understanding tasks, especially in situations involving multimodal information and cultural backgrounds.
@@
To address these problems, inspired by how humans (possibly) understand metaphors, we find that the essence of the difficulty in metaphor understanding and generation is contextual missing. Therefore, we propose a novel framework that more closely aligns with human cognitive processes for metaphor interpretation. Our framework first transforms visual information into textual representations and then iteratively searches to enrich these representations with out-of-domain knowledge, enabling deeper inferential reasoning. Experiments from both Multiple-Choice Question and Open-Style Question consistently verify the superiority of the proposed framework.
@@
  • We systematically analyze image implication tasks and find the difficulty of the image implication understanding and reasoning task lies in contextual missing. From the perspective of human cognition, we proposed a new direction for solving these tasks – Contextual Alignment.
  • We propose a novel human-inspired three-stage framework Let Androids Dream (LAD) for image implication understanding and reasoning, including Perception, Search and Reasoning. Our LAD implements the lightweight GPT-4o-mini model to achieve SOTA on English image implication benchmark (1300+ questions) and a huge improvement on Chinese image implication benchmark (800 questions), comparable with the Gemini-3.0-pro model and other top closed-source models on Multiple-Choice Question (MCQ). Generalization experiments also show that our framework can effectively benefit general VQA and visual reasoning tasks.
  • We design the challenging Open-Style Question (OSQ) with comprehensive metric to automatically evaluate the image implication tasks. This metric aligns 95.7% with human annotations, making it more suitable for diverse evaluation. Our LAD outperforms the GPT-4o model 36.7% on OSQ.
@@ 2.1 Image Implication
Image implication encompasses various cognitive aspects, including humor, sarcasm, and broader metaphorical understanding. Early research in this domain focused on specialized aspects, such as humor recognition [hessel-etal-2023-androids, horvitz-etal-2024-getting] and sarcasm detection [Desai_Chakraborty_Akhtar_2022]. As the rapid development of large language models (LLMs) brings new opportunities for analyzing image implication, we need more comprehensive evaluation frameworks. DeepEval [yang-etal-2024-large] provided a systematic taxonomy of image implications. Subsequently, II-Bench [liu2024iibenchimageimplicationunderstanding] emerged as the first English image implication benchmark, followed by CII-Bench [zhang2024mllmsunderstanddeepimplication], which extended this evaluation framework to Chinese images. Image implication understanding requires sophisticated multi-hop reasoning and theory of mind (ToM) capabilities [liu2024iibenchimageimplicationunderstanding, zhang2024mllmsunderstanddeepimplication]. Existing approaches fall into two categories: explicit metaphor mapping and model implicit reasoning. The first approach, represented by CLOT [zhong2024lets], constructs mappings between metaphor ontologies and visual representations. However, this approach faces key challenges: metaphorical relationships have complex many-to-many mappings that are difficult to formalize, and cultural references are too dynamic for static mappings. The second approach, exemplified by C4MMD [xu-etal-2024-exploring], employs training-free CoT reasoning. Despite its promise, this approach struggles with the complex nature of metaphorical understanding, which surpasses traditional reasoning. The large search space for out-of-domain reasoning and changing cultural contexts limits its effectiveness. To address this, we propose a novel methodology that transforms visual information into texts and iteratively enriches them with out-of-domain knowledge, better aligning with human cognitive processes.
@@ 2.2 Vision-language Reasoning
The rapid advancement of LLMs has demonstrated remarkable text reasoning capabilities, as evidenced by models such as o1 [o1], DeepSeek-R1 [deepseek-R1], and QwQ [qwq-32b-preview, qwen2]. However, real-world knowledge often transcends textual representation, with visual information encapsulating substantial world knowledge that pure language models cannot access. For example, images inherently contain rich, multi-layered information that often resists straightforward textual description, including spatial relationships, contextual nuances, and implicit knowledge that humans process intuitively. This limitation has driven research toward integrating visual information into text-based reasoning frameworks. Current research has developed three primary approaches to incorporate visual information into model reasoning: 1) Comprehensive MLLM Description: This approach treats visual content as a text grounding problem, as demonstrated by LLAVA-COT [xu2024llavacotletvisionlanguage] and Mulberry [yao2024mulberryempoweringmllmo1like]. 2) Multi-turn MLLM Interaction: Models like VoCoT [li2024vocotunleashingvisuallygrounded] and V* [vstar] employ iterative question-answering to extract fine-grained visual information at various levels of detail. 3) Tool-augmented Reasoning: Frameworks such as Visual Sketchpad [hu2024visual] and Whiteboard-of-Thought [menon2024whiteboard] leverage tool-based approaches to modify images and augment reasoning with prior knowledge embedded in these tools.
@@ 3.2 Stage II: Search
    WebSearch: For questions requiring external, dynamic, or highly specific information, WebSearch is invoked. Inspired by LLM search methods like MindSearch [chen2024mindsearch], but focusing on image implication tasks, our WebSearch component first employs the planner. The planner, acting as a high-level strategist, decomposes the initial search question into a series of more granular sub-questions. These sub-questions are structured into a directed acyclic graph (DAG), simulating a multi-step, exploratory information-seeking process. Subsequently, the searcher executes this plan. It performs hierarchical information retrieval for each sub-question from the internet, gathering relevant snippets and facts. This multi-agent method, with distinct planner and searcher modules, allows for parallel processing and dynamic refinement of the search strategy. The retrieved information for sub-questions is then synthesized to answer the original search question. This ensures access to recent developments and a broad spectrum of public knowledge, crucial for understanding contemporary image implications.
@@ 4.1 Baselines
Qwen2.5-VL-7B [qwen25vl] 46% 40% 2.34 2.58
DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 46% 36% 2.82 2.86
GLM-4.1V-8B [glm4v] 60% 52% 2.60 2.96
Gemini-2.0-flash [gemini] 70% 68% 1.60 3.12
Qwen2.5-VL-72B [qwen25vl] 72% 56% 1.56 3.12
InternVL3-78B [internvl3] 70% 74% 3.42 3.70
GLM-4V-plus [glm4v] 64% 64% 3.01 3.12
Grok-3 [grok3] 66% 64% 3.24 2.96
Claude-3.5-Sonnet [claude35sonnet] 68% 62% 3.22 3.78
GPT-4o [gpt4o] 74% 58% 2.94 3.76
GPT-4.1 [gpt4o] 74% 62% 3.30 3.92
@@
Gemini-2.0-flash-thinking [gemini] 64% 68% 1.66 2.84
QVQ-72B [qvq-72b-preview] 62% 56% 3.10 3.42
Doubao-1.5-thinking-vision-pro [doubao-1.5-thinking-vision-pro] 66% 66% 3.16 3.90
Grok-3-reasoning [grok3] 74% 64% 3.06 2.92
Gemini-3.0-pro [gemini3] 76% 76% 3.82 3.96
@@
Table 1: Overall results of different models on Multiple-Choice Question and Open-Style Question. The best-performing model in each category is in-bold, and the second best is underlined. Performance differences relative to base models are shown as colodarkred subscripts: for improvements, for declines.
@@
Evaluation. Our evaluation utilizes two comprehensive image implication benchmarks, II-Bench [liu2024iibenchimageimplicationunderstanding] and CII-Bench [zhang2024mllmsunderstanddeepimplication], both featuring Multiple-Choice Question (MCQ). Furthermore, we manually construct the high-level benchmark by selecting 100 high-quality, diverse and representative images from varied image types like illustrations and comics. The detailed statistic is in Appendix D. And we measure accuracy by comparing the model’s selected option to the ground truth. Aware of potential MCQ biases [li2024multiplechoicequestionsreallyuseful, OpenLLMLeaderboard, zheng2024largelanguagemodelsrobust] and the greater difficulty of generation over judgment tasks, we introduce a novel evaluation method Open-Style Question (OSQ). It uses the same images with the fixed question: “What is the implication in this image?”. And we use GPT-4o with a specialized evaluation metric as evaluators, validated by multiple human consistency checks. We also conduct a further analysis of experiments’ findings in Appendix E.
Figure 3: A case study of different methods on Multiple-Choice Question. The End2End method shows superficial reasoning and the CoT method shows over-inference, while our LAD framework shows the correct path via more contextual alignment analysis. The full prompt is listed in Appendix F.
@@ 4.2.2 Results and Analysis
Table 1 presents comprehensive results of MCQ across different MLLMs on our high-level benchmark. The LAD framework demonstrates remarkable effectiveness, achieving SOTA performance with the lightweight GPT-4o-mini model. In English MCQ, our framework matches the performance of top closed-sourced model Gemini-3.0-pro, while significantly outperforming Claude-3.5-Sonnet by 9%. For Chinese MCQ, our framework achieves comparable results to GPT-4o, while substantially surpassing DeepSeek-VL2 by 44.4%.